A firm has an EPS of $2.50, a book value per share of $23.50, and a market/book ratio of 2.7x. What is the P/E Ratio?
Answer #1

Market/book ratio = Market price per share / Book value per share

2.7 = Market price per share / $23.50

Market price per share = 2.7 * $23.50

Market price per share = $63.45

P/E Ratio = Market price per share / EPS

P/E Ratio = $63.45 / $2.50

P/E Ratio = 25.38